# The Astrometric Resoeccentric Degeneracy: Eccentric Single Planets Mimic 2:1 Resonant Planet Pairs in Astrometry

### v1

@@ Abstract
software: This work made use of the following software packages: matplotlib (Hunter:2007), numpy (numpy), python (python), JAX (jax2018github), jaxoplanet (jaxoplanet), and gaiascanlaw. Software citation information aggregated using The Software Citation Station (software-citation-station-paper; software-citation-station-zenodo).
1 Introduction
@@
Astrometric detection of exoplanets is of particular interest at the moment as we approach the era of Gaia epoch astrometry (Gaia2016). Gaia Data Release 4 (DR4) is scheduled to release the first $\sim$5.5 years of time series astrometry for up to $\sim$2.7 billion sources across the complete sky in December 2026.111According to https://www.cosmos.esa.int/web/gaia/data-release-4 accessed on November 30, 2025. Gaia’s ability to search for giant exoplanets on distant orbits via the reflex motion they induce on their host stars has long been appreciated (Bernstein1995; Casertano1996; Lattanzi2000; Casertano2008; Perryman2014; Sozzetti2014; Ranalli2018). Recently, lammers2025exoplanet re-investigated the exoplanet yield from Gaia DR4 and the subsequent mission-length $\sim$10.5 years of time series astrometry in Gaia DR5. This analysis included improvements in our understanding of distant giant-planet occurrence, the local stellar population, and Gaia astrometric precision. Via injection recovery tests, lammers2025exoplanet demonstrated that Gaia will be most sensitive to super-Jupiters ($3-13\,M_{\rm Jup}$) on $2-5\,\mathrm{AU}$ orbits around GKM-type stars ($0.4-1.3\,M_{\odot}$) within $500\,\mathrm{pc}$, predicting $\sim$1,900 planets in DR4 and $\sim$38,000 planets in DR5 with masses and orbital periods determined to better than 20%. However, they assumed an underlying population solely composed of single-planet systems, which, of course, will not be the true underlying population. An important scientific result from Gaia will be precise constraints on cool gas giant occurrence rates and so it is essential to understand degeneracies and multi-modalities that will be present in the Gaia DR4 and DR5 datasets in order to not bias these occurrence rate determinations.
@@
Radial-velocity (RV) surveys are known to suffer from a degeneracy in which a system of two planets on nearly circular orbits in or near a 2:1 mean-motion resonance can be well fit by a single-planet model with a moderately eccentric orbit – henceforth called the resoeccentric degeneracy. This effect was first analyzed in detail by AngladaEscude2010, where it was shown that $\sim$35% of published single-planet eccentric solutions are statistically compatible with circular two-planet 2:1 configurations. In the RV case, the Keplerian signal of a single eccentric orbit can be expanded as a sum of a fundamental mode at frequency $n$ and a first harmonic at frequency $2n$ – with the harmonic amplitude $\simeq e$ times the fundamental. A two-planet circular 2:1 configuration naturally reproduces the same harmonic structure, as the outer planet mimics the fundamental frequency and the inner planet mimics the first harmonic.
@@
The connection between this degeneracy and astrometric observations was predicted in AngladaEscude2010 – where it was noted that astrometry could provide ways to uncover eccentric imposters for systems in non-coplanar orientations where the inherent degeneracies will be broken. However, the analytics were not explicitly presented for the astrometric case; rather, AngladaEscude2010 points the reader to analyses of frequency decomposition for astrometric detection in Konacki2002.
We have organized this paper as follows. In §2 we derive the astrometric degeneracy between circular coplanar 2:1 resonant systems and single eccentric orbits via fundamental and first harmonic frequency analysis (good to first order in eccentricity). In §3 we demonstrate this degeneracy via simulated Gaia DR4 ($\sim$5.5 years) and DR5 ($\sim$10.5 years) data. In §4 we discuss the implications of this degeneracy in astrometric datasets, such as Gaia, and its impact on interpreting the demographics that arise from Gaia as well as the dynamical implications of potential detections of mutually inclined 2:1 resonant planets.
2 The Resoeccentric Degeneracy
Here we derive the astrometric resoeccentric degeneracy. In this derivation, we adopt the orbital coordinate frame as described in Feng2024 and assume purely Keplerian orbits.
The apparent astrometric displacement can be described in terms of the scaled Thiele-Innes constants $A^{\prime},B^{\prime},F^{\prime},G^{\prime}$ (Thiele1883), defined as
@@ 2 The Resoeccentric Degeneracy
2.1 Single mildly eccentric orbit: harmonic structure
@@
2.2 Two circular planets in a 2:1 coplanar configuration
Now consider two planets on circular orbits around the same star with periods $P_{1}=2P_{2}$ and stellar reflex semi-major axes $a_{\star,1}$ and $a_{\star,2}$. We assume that the orbits are coplanar and share the same $(\Omega,i,\omega)$, so that the same scaled Thiele-Innes constants $(A^{\prime},B^{\prime},F^{\prime},G^{\prime})$ apply to both planets. Assuming the outer planet shares an orbital period with the single planet analyzed in §2.1 so $P_{1}=P$, in the orbital plane the total displacement may be written as a linear combination of contributions from each planet to zeroth order in eccentricity, as we assume circular orbits, or $e_{1}=e_{2}=0$:
@@ 2.2 Two circular planets in a 2:1 coplanar configuration
We note that systems in exact mean motion resonant configurations will be significantly influenced by N-body interactions, causing additional effects that may differentiate the signals from Keplerian orbits (beague2013resonance; batygin2013analytic). For systems in near 2:1 mean motion resonance – a common detected configuration in exoplanet architectures (e.g., Fabrycky2014) – this Keplerian approximation should hold (MurrayDermott1999).
2.3 Effective eccentricity mapping
@@
where $e_{\rm eff}$ is the effective eccentricity of the single-planet model. Solving for $e_{\rm eff}$,
@@
In other words, the ratio of the astrometric semi-major axes induced by the inner and outer planets sets the effective eccentricity of the single-planet orbit that reproduces the same time-dependent harmonic content (up to first order in $e$ and ignoring the constant offset). Because the Thiele-Innes constants for the two planets are identical in the coplanar, aligned case, this mapping holds simultaneously in both coordinates.
@@
2.4 From harmonic amplitudes to effective eccentricity
The astrometric resoeccentric degeneracy can be related to physical quantities of interest by converting the effective eccentricity $e_{\rm eff}$ into a relation involving the planet masses and periods. The astrometric semi-major axis induced by planet $k$ for a circular orbit is (Sozzetti2005)
@@ 2.4 From harmonic amplitudes to effective eccentricity
Equation (20) shows that the apparent eccentricity of a single-planet astrometric fit is not set solely by the inner planet’s mass relative to the outer planet, but also by the geometric lever arm provided by the period (and therefore semi-major axis) ratio. For a strict 2:1 configuration with $P_{2}=P_{1}/2$, we have
@@
For instance, in this exact 2:1 case, an apparent eccentricity of $e_{\rm eff}=0.30$ corresponds to $M_{p,2}/M_{p,1}\approx 0.24$, meaning that an inner planet with only one quarter of the mass of the outer planet can mimic the harmonic signature of a single planet with $e\approx 0.3$. Thus, the astrometric resoeccentric degeneracy maps cleanly into a mass-period relation, providing a physically intuitive explanation of how an apparently eccentric orbit can arise from two circular, phase-aligned planets in resonance.
2.5 Mutual inclination breaks the degeneracy
@@
Consequently, for circular-orbiting planets the degeneracy is broken only by differences in inclination or longitude of ascending node. Mutual inclination produces distinct Thiele-Innes constants and therefore destroys the resoeccentric degeneracy. In contrast, perfectly coplanar circular planets with the same $(i,\Omega)$ remain degenerate with a single eccentric orbit at first order, regardless of differences in $\omega$ or in orbital phase.
3 Simulations
We demonstrate how Gaia astrometric observations may have difficulties in differentiating between two-planet systems in a 2:1 mean-motion commensurability and a single eccentric planet. Using the jaxoplanet (jaxoplanet) framework, we create a Keplerian orbit for a pair of $e=0$ giant planets with masses $M_{p,1}=12\,M_{\mathrm{Jup}}$ and $M_{p,2}=(0.79)\,(e_{\rm eff}=0.3)\,({M_{p,1}})=2.84\,M_{\mathrm{Jup}}$ and with orbital periods $P_{1}$ and $P_{2}=P_{1}/2$. Following the same procedure, we additionally create a Keplerian orbit for a single eccentric planet of mass $M_{p,1}=12\,M_{\mathrm{Jup}}$, period $P_{1}$, and eccentricity $e_{\mathrm{eff}}=0.3$. We then duplicate this effort for a system with a mutual inclination of $45^{\circ}$ to demonstrate the resoeccentric degeneracy breaking capabilities of mutually inclined orbits.
@@ 3 Simulations
Gaia does not measure instantaneous RA and Dec, but rather the along–scan (AL) coordinate $\Delta\eta$ at each field-of-view transit. For every simulated epoch $t_{k}$ with scan angle $\psi_{k}$ (computed from the full Gaia scanning law using gaiascanlaw), we project the sky-plane reflex motion into the AL direction via
@@
With this approach, we can generate synthetic sky–plane (RA, Dec) as well as along-scan ($\Delta\eta$) measurements for Gaia DR4 (5.5 yrs) and Gaia DR5 (10.5 yrs). The resulting sky-plane orbits, RA/Dec time series for these simulations, and along-scan astrometric measurements ($\Delta\eta)$ are shown in Figure 1 and Figure 2. These demonstrate that [1] it is possible for a single eccentric planet to mimic a coplanar, circular 2:1 resonant planetary system in both Gaia DR4 and DR5 and [2] in certain orbital & mass configurations, mutual inclination will break this degeneracy – allowing observers to statistically distinguish between these two model hypotheses. Future work should investigate for which orbital configurations precisely will this degeneracy breaking occur in Gaia DR4 and DR5.
Figure 1: Gaia DR5-like simulated orbits (vertical line shows Gaia DR4 baseline), coplanar system ($T_{\mathrm{span}}=5.5\,\mathrm{yr}$, $i_{\mathrm{mut}}=0^{\circ}$): 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single-planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pc. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star; $\Delta\eta$ is the Gaia along–scan (AL) coordinate; Res are the residuals between the simulated Gaia-like data with the eccentric model and the 2:1 model, respectively.
@@
Figure 2: Gaia DR5-like simulated orbits (vertical line shows Gaia DR4 baseline), mutually inclined system ($T_{\mathrm{span}}=5.5\,\mathrm{yr}$, $i_{\mathrm{mut}}\approx 45^{\circ}$): 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single-planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pcpc. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star; $\Delta\eta$ is the Gaia along–scan (AL) coordinate; Res are the residuals between the simulated Gaia-like data with the eccentric model and the 2:1 model, respectively.
4 Implications
4.1 Exoplanet Astrometry in the Gaia Era

### v2

@@ Abstract
software: This work made use of the following software packages: matplotlib (Hunter, 2007), numpy (Harris et al., 2020), python (Van Rossum and Drake, 2009), JAX (Bradbury et al., 2018), jaxoplanet (Hattori et al., 2024), and gaiascanlaw (Penoyre, 2025) Software citation information aggregated using The Software Citation Station (Wagg and Broekgaarden, 2024; Wagg et al., 2025).
I Introduction
@@
Astrometric detection of exoplanets is of particular interest at the moment as we approach the era of Gaia epoch astrometry (Gaia Collaboration et al., 2016). Gaia Data Release 4 (DR4) is scheduled to release the first $\sim$5.5 years of time series astrometry for up to $\sim$2.7 billion sources across the complete sky in December 2026.111According to https://www.cosmos.esa.int/web/gaia/data-release-4 accessed on November 30, 2025. Gaia’s ability to search for giant exoplanets on distant orbits via the reflex motion they induce on their host stars has long been appreciated (Bernstein and Bastian, 1995; Casertano et al., 1996; Lattanzi et al., 2000; Casertano et al., 2008; Perryman et al., 2014; Sozzetti et al., 2014; Ranalli et al., 2018). Recently, Lammers and Winn (2025) re-investigated the exoplanet yield from Gaia DR4 and the subsequent mission-length $\sim$10.5 years of time series astrometry in Gaia DR5. This analysis included improvements in our understanding of distant giant-planet occurrence, the local stellar population, and Gaia astrometric precision. Via injection recovery tests, Lammers and Winn (2025) demonstrated that Gaia will be most sensitive to super-Jupiters ($3-13\,M_{\rm Jup}$) on $2-5\,\mathrm{AU}$ orbits around GKM-type stars ($0.4-1.3\,M_{\odot}$) within $500\,\mathrm{pc}$, predicting $\sim$1,900 planets in DR4 and $\sim$38,000 planets in DR5 with masses and orbital periods determined to better than 20%. However, they assumed an underlying population solely composed of single planet systems, which, of course, will not be the true underlying population. An important scientific result from Gaia will be precise constraints on cool gas giant occurrence rates and so it is essential to understand degeneracies and multi-modalities that will be present in the Gaia DR4 and DR5 datasets in order to not bias these occurrence rate determinations.
@@
Radial-velocity (RV) surveys are known to suffer from a degeneracy in which a system of two planets on nearly circular orbits in or near a 2:1 mean-motion resonance can be well fit by a single planet model with a moderately eccentric orbit – henceforth called the resoeccentric degeneracy. This effect was first analyzed in detail by Anglada-Escudé et al. (2010), where it was shown that $\sim$35% of published single planet eccentric solutions are statistically compatible with circular two-planet 2:1 configurations. A subsequent Bayesian model comparison study on a more selective sample of RV eccentric single systems found similar results – showing that 25% of the 60 selected systems are likely to be mischaracterized and are in fact two planet systems at a 95% confidence level (Boisvert et al., 2018). In the RV case, the Keplerian signal of a single eccentric orbit can be expanded as a sum of a fundamental mode at frequency $n$ and a first harmonic at frequency $2n$ – with the harmonic amplitude $\simeq e$ times the fundamental. A two-planet circular 2:1 configuration naturally reproduces the same harmonic structure, as the outer planet mimics the fundamental frequency and the inner planet mimics the first harmonic.
@@
The connection between this degeneracy and astrometric observations was predicted in Anglada-Escudé et al. (2010) – where it was noted that astrometry could provide ways to uncover eccentric imposters for systems in non-coplanar orientations where the inherent degeneracies will be broken. However, the mathematical derivation was not explicitly presented for the astrometric case; rather, Anglada-Escudé et al. (2010) points the reader to analyses of frequency decomposition for astrometric detection in Konacki et al. (2002).
We have organized this paper as follows. In §II we derive the astrometric degeneracy between circular coplanar 2:1 resonant systems and single eccentric orbits via fundamental and first harmonic frequency analysis (good to first order in eccentricity). In §III we demonstrate this degeneracy via simulated Gaia DR4 ($\sim$5.5 years) and DR5 ($\sim$10.5 years) data. In §IV we discuss the implications of this degeneracy in astrometric datasets, such as Gaia, and its impact on interpreting the demographics that arise from Gaia as well as the dynamical implications of potential detections of mutually inclined 2:1 resonant planets.
II The Resoeccentric Degeneracy
Here we derive the astrometric resoeccentric degeneracy. In this derivation, we adopt the orbital coordinate frame as described in Feng (2024) and assume purely Keplerian orbits.
The apparent astrometric displacement can be described in terms of the scaled Thiele-Innes constants $A^{\prime},B^{\prime},F^{\prime},G^{\prime}$ (Thiele, 1883), defined as
@@ II The Resoeccentric Degeneracy
II.1 Single moderately eccentric orbit: harmonic structure
@@
II.2 Two circular planets in a 2:1 coplanar configuration
Now consider two planets on circular orbits around the same star with periods $P_{1}=2P_{2}$ and stellar reflex semi-major axes $a_{\star,1}$ and $a_{\star,2}$. We assume that the orbits are coplanar and share the same $(\Omega,i,\omega)$, so that the same scaled Thiele-Innes constants $(A^{\prime},B^{\prime},F^{\prime},G^{\prime})$ apply to both planets. Assuming the outer planet shares an orbital period with the single planet analyzed in §II.1 so $P_{1}=P$, in the orbital plane the total displacement may be written as a linear combination of contributions from each planet to zeroth order in eccentricity, as we assume circular orbits, or $e_{1}=e_{2}=0$:
@@ II.2 Two circular planets in a 2:1 coplanar configuration
We note that systems in exact mean motion resonant configurations will be influenced by N-body interactions, causing additional effects that may differentiate the signals from Keplerian orbits (Beaugé et al., 2003; Batygin and Morbidelli, 2013a). For systems in near 2:1 mean motion resonance – a common detected configuration in exoplanet architectures (e.g., Fabrycky et al., 2014) – this Keplerian approximation should hold (Murray and Dermott, 1999).
II.3 Effective eccentricity mapping
@@
where $e_{\rm eff}$ is the effective eccentricity of the single planet model. Solving for $e_{\rm eff}$,
@@
In other words, the ratio of the astrometric semi-major axes induced by the inner and outer planets sets the effective eccentricity of the single planet orbit that reproduces the same time-dependent harmonic content (up to first order in $e$ and ignoring the constant offset). Because the Thiele-Innes constants for the two planets are identical in the coplanar, aligned case, this mapping holds simultaneously in both coordinates.
@@
II.4 From harmonic amplitudes to effective eccentricity
The astrometric resoeccentric degeneracy can be related to physical quantities of interest by converting the effective eccentricity $e_{\rm eff}$ into a relation involving the planet masses and periods. The astrometric semi-major axis induced by planet $k$ for a circular orbit is (Sozzetti, 2005)
@@ II.4 From harmonic amplitudes to effective eccentricity
Equation (20) shows that the apparent eccentricity of a single planet astrometric fit is not set solely by the inner planet’s mass relative to the outer planet, but also by the geometric lever arm provided by the period (and therefore semi-major axis) ratio. For a strict 2:1 configuration with $P_{2}=P_{1}/2$, we have
@@
For instance, in this exact 2:1 case, an apparent eccentricity of $e_{\rm eff}=0.30$ corresponds to $M_{p,2}/M_{p,1}\approx 0.24$, meaning that an inner planet with only one quarter of the mass of the outer planet can mimic the harmonic signature of a single planet with $e\approx 0.3$. Thus, the astrometric resoeccentric degeneracy maps cleanly into a mass-period relation, providing a physically intuitive explanation of how an apparently eccentric orbit can arise from two circular, phase-aligned planets in resonance. As eccentricity will always be less than unity, in order for two planets to mimic a single eccentric planet their mass ratio must be less than $\sim$0.79 with the inner planet having the smaller mass of the two planets.
II.5 Mutual inclination breaks the degeneracy
@@
Consequently, for circular-orbiting planets the degeneracy is broken only by relative differences between the two planets’ inclinations or longitudes of ascending node. Mutual inclination or nodal offsets produce distinct Thiele-Innes constants, therefore resolving the resoeccentric degeneracy. In contrast, perfectly coplanar circular planets with the same $(i,\Omega)$ remain degenerate with a single eccentric orbit at first order, regardless of differences in $\omega$ or in orbital phase.
The precise mutual inclination at which this degeneracy breaking will occur will depend on a number of factors – including the masses and periods of the planets in the system, how far from a perfect 2:1 configuration are the two planets, the mass of the star, the distance to the star, and the location of the star in the sky (which affects the cadence and number of Gaia observations) as well as modeling and sampling choices. As such, we leave a detailed analysis of this to future work.
III Simulations
Figure 1: Gaia-like simulated orbits (vertical line shows Gaia DR4 baseline), coplanar system ($T_{\mathrm{span}}=5.5\,\mathrm{yr}$, $i_{\mathrm{mut}}=0^{\circ}$): 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pc. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star; $\Delta\eta$ is the Gaia along–scan (AL) coordinate; Res are the residuals between the simulated Gaia-like data with the eccentric model and the 2:1 model, respectively.
We demonstrate how Gaia astrometric observations may have difficulties in differentiating between two-planet systems in a 2:1 mean-motion commensurability and a single eccentric planet. Using the jaxoplanet (Hattori et al., 2024) framework, we create a Keplerian orbit for a pair of $e=0$ giant planets with masses $M_{p,1}=12\,M_{\mathrm{Jup}}$ and $M_{p,2}=(0.79)\,(e_{\rm eff}=0.3)\,({M_{p,1}})=2.84\,M_{\mathrm{Jup}}$ and with orbital periods $P_{1}$ and $P_{2}=P_{1}/2$. Following the same procedure, we additionally create a Keplerian orbit for a single eccentric planet of mass $M_{p,1}=12\,M_{\mathrm{Jup}}$, period $P_{1}$, and eccentricity $e_{\mathrm{eff}}=0.3$. We then duplicate this effort for a 2:1 system with a range of mutual inclinations set to $\{0,15,30,45\}^{\circ}$ to demonstrate the resoeccentric degeneracy breaking capabilities of mutually inclined orbits.
@@ III Simulations
Gaia does not measure instantaneous RA and Dec, but rather the along–scan (AL) coordinate $\Delta\eta$ at each field-of-view transit. For every simulated epoch $t_{k}$ with scan angle $\psi_{k}$ (computed from the full Gaia scanning law using gaiascanlaw (Penoyre, 2025)), we project the sky-plane reflex motion into the AL direction via
@@
Figure 2: Gaia-like simulated orbits showing that at some mutual inclination the astrometric resoeccentric degeneracy breaks. We simulate the same systems as in Figure 1, but with mutual inclinations of $i_{\mathrm{mut}}=\{0,15,30,45\}^{\circ}$ for a 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pcpc [left] or 100 pcpc [right]. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star.
With this approach, we can generate synthetic sky–plane (RA, Dec) reflex motion as well as along-scan ($\Delta\eta$) measurements for Gaia DR4 (5.5 yrs) and Gaia DR5 (10.5 yrs). The resulting sky-plane orbits, RA/Dec time series for these simulations, and along-scan astrometric measurements ($\Delta\eta)$ are shown in Figure 1. In Figure 2, we reproduce this effort showing only the sky–plane (RA, Dec) for a range of mutual inclinations ($i_{\mathrm{mut}}=\{0,15,30,45\}^{\circ}$) for the 2:1 system and the host star at 50 $pc$ and 100 $pc$. These demonstrate that [1] it is possible for a single eccentric planet to mimic a coplanar, circular 2:1 resonant planetary system in both Gaia DR4 and DR5 and [2] in certain orbital & mass configurations, mutual inclination will break this degeneracy – allowing observers to statistically distinguish between these two model hypotheses. We emphasize that these figures are only for a single planetary orbital-mass configuration and should not be taken as representative of the full population of Gaia DR4 and DR5 planetary systems – but rather a proof-of-concept for Gaia-like data. Future work should investigate for which orbital configurations precisely will this degeneracy breaking occur in Gaia DR4 and DR5.
IV Implications
The astrometric resoeccentric degeneracy is a mathematical degeneracy that is perfect to first order in eccentricity and under the assumption of “perfect data.” In this section, we discuss some of the consequences of this degeneracy in the context of the imminent release of Gaia DR4, as well as the dynamical histories implied by planets in 2:1 MMR configurations. While our discussions regarding the likely architectures of these 2:1 MMR systems rest upon an extensive legacy of past work, we briefly opine that Gaia DR4 and DR5 will significantly improve our population-level understanding of this range of exoplanetary orbital-mass parameter space: just as other pioneering discoveries (e.g., Mayor and Queloz, 1995) inspired revisions of theoretical outcomes of planet formation, the astrometric exoplanet revolution may similarly reshape current expectations. In this context, there is value in remaining open to a broader range of possible planetary architectures, recognizing that those emphasized in the literature largely reflect current observational capabilities.
IV.1 Exoplanet Astrometry in the Gaia Era